At St. Eligius Hospital, the weights of newborn babies follow a Normal distribution, with a mean of 7.5 pounds and a standard deviation of 1.2 pounds. If Dr. Cavanero tells a mother that her newborn baby has a weight that is at the third quartile (or Q3), this means the baby must weigh approximately A. 6.66 pounds. B. 7.75 pounds. C. 8.34 pounds. D. 8.75 pounds. E. 9.05 pounds
Answer #1

P(Z < z) = 75%

= P(Z < z) = 0.75  

= P(Z < 0.6745 ) = 0.75

z = 0.6745

Using z-score formula,

x = z * \sigma + \mu

x = 0.6745 * 1.2+7.5

x = 8.34